Electric Field Modelling in an Electrostatic Intrabody Signalling

System by Matthew Gray

The performance of data transmission by electrostatic signalling through the human body is highly subject to the geometry of the system. The system is modelled as a set of equipotentials, with their capacitive couplings solved for based on the current geometry. This paper describes various transmitter/receiver geometries and presents theoretical calculations for relative signal strengths.